OpenStudy (anonymous):

Solve: The quantity 2 x minus 20 divided by 3 = 2x

Vocaloid (vocaloid):

(2x-20)/3 = 2x start by multiplying both sides by 3

OpenStudy (anonymous):

by both sides do you mean .. both sides of the equation ?

Vocaloid (vocaloid):

yes

OpenStudy (usukidoll):

\[\large \frac{2x-20}{3} =2x \] yes multiply 3 throughout the equation

OpenStudy (jdoe0001):

\(\bf {\color{brown}{ \cancel{3}}}\cdot \cfrac{2x-20}{\cancel{3}} ={\color{brown}{ 3}}\cdot 2x\)

OpenStudy (anonymous):

so would i have 6x-60=6x ?

OpenStudy (usukidoll):

by multiplying 3 on both sides... the 3 for the left side ends up being canceled out

OpenStudy (anonymous):

so .... 2x-20= 3 times 2x ? im lost .. im not good at math

OpenStudy (usukidoll):

or another way of seeing that there is cancellation \[\large \frac{(3)(2x-20)}{3} =2x (3)\] \[\large \frac{(6x-60)}{3} =2x (3)\] \[\large 2x-20=6x\]

OpenStudy (usukidoll):

yeah you will have 2x-20= 3(2x) which becomes 2x-20=6x

OpenStudy (anonymous):

okay i understand that part now

OpenStudy (usukidoll):

now we just have to have all variables to the left and all numbers to the right so add 20 to both sides and subtract 6x to both sides

OpenStudy (anonymous):

can i solve it like a regular equation ....